6. Repeat Problem 5, but use constant specific heats at the average temperature of (900 K+310 K)2. This will give results tha5. A gas-turbine power plant operates on the simple Brayton cycle with air as the working fluid and delivers 32 MW of power.
Answer #1

© 1-5 diagoon T- 90 kn Qout. 310k t =) Given pressure ratio op = 8 900 + 310 = 605 K Targ = 2 2 from aig tables At 605k ep =Ne = . wrev wach ofc T2-T) CP (T2-TI) 0.80 54 3.57 - 310 T2 - 310 Tz - 610.96K. IP? 900 T S 513:26K. 1.371 (8) 1-37 = = n APower output = 32 mw = 32x103 kw a power = micp (B3-T4) 32000 = mix 1.053 ( 900- 567.40] m = 91.36 kgs. mass = ကံ ၃ ပT = 91.3